# Groupoid Graded Semisimple Rings

### v1

@@ 4.2. Graded modules over gr-division rings
The proofs of Theorem 4.4 and Corollary 4.6 follow very much the pattern for group graded division rings [17, Section 1.4].
@@
Let $D$ be a $\Gamma$-graded division ring and $M$ be a $\Gamma$-graded $D$-module. The following assertions hold true.
@@
[…]
@@
If $D$ is a $\Gamma$-graded division ring, then every $\Gamma$-graded $D$-module has a pseudo-basis by Theorem 4.4(1). Unlike in the group graded context [4, Theorem 3.3], the converse is not true. We will deal with this situation in Section 7.
@@
Let $D$ be a $\Gamma$-graded division ring and $M$, $N$ be $\Gamma$-graded $D$-modules. For all $(\gamma_{i})_{i\in I}\in\Gamma^{I}$, pseudo-linear independent sequence $(x_{i})_{i\in I}\in\prod\limits_{i\in I}M_{\gamma_{i}}$, $\sigma\in\Gamma$ and $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N_{\sigma\gamma_{i}}$, there exists $g\in\operatorname{HOM}_{D}(M,N)_{\sigma}$ such that $g(x_{i})=y_{i}$ for each $i\in I$.
@@ 4.3. Graded linear algebra over gr-division rings
$\operatorname{M}_{m\times n}(R)[\overline{\alpha}][\overline{\beta}]=\left\{\begin{pmatrix}a_{11}&\cdots&a_{1n}\\ \vdots&\ddots&\vdots\\ a_{m1}&\cdots&a_{mn}\end{pmatrix}\in\operatorname{M}_{m\times n}(R)\ \bigg{|}\ a_{ij}\in R_{\alpha_{i}\beta_{j}^{-1}}\right\}.$
@@
Define $[f]_{\mathcal{B}\mathcal{C}}\in\operatorname{M}_{m\times n}(R)[\overline{\alpha}][\overline{\beta}]$ as the matrix
$[f]_{\mathcal{BC}}=\Big{[}(f(u_{1}))_{\mathcal{C}}\ (f(u_{2}))_{\mathcal{C}}\ \dotsb(f(u_{n}))_{\mathcal{C}}\Big{]}=\begin{pmatrix}a_{11}&a_{12}&\ldots&a_{1n}\\ a_{21}&a_{22}&\ldots&a_{2n}\\ \vdots&\ddots&\ddots&\vdots\\ a_{m1}&a_{m2}&\ldots&a_{mn}\end{pmatrix}.$
@@
$[f]_{\mathcal{BC}}\cdot(x)_{\mathcal{B}}=(f(x))_{\mathcal{C}}\ \textrm{ for all }x\in\operatorname{h}(M)$ (4.1)
and that $[f]_{\mathcal{BC}}$ is the unique matrix in $\operatorname{M}_{m\times n}(R)[\overline{\alpha}][\overline{\beta}]$ that satisfies (4.1).
@@
$[f]_{\mathcal{BC}}\cdot[g]_{\mathcal{AB}}=[f\circ g]_{\mathcal{AC}}.$
@@
The following result, shows that all for ranks just defined are equal over a graded division ring. We follow very close the proof in [23, Exercises 13.13, 13.14] where the result is proved for (ungraded) division rings.
@@ 5.1. Gr-simple modules and Schur’s lemma
When considering rings and modules without a grading, it is useful to study simple modules up to isomorphism. In the graded context, one has to take into account the shifts of the simple modules. Inspired by the idea of [10, p. 395], we say that two $\Gamma$-graded $R$-modules $M$ and $N$ are in the same isoshift class if there exists $\Sigma\subseteq\Gamma$ such that $\Sigma$ is fully $r$-unique for $N$, $\Sigma^{-1}$ is fully $r$-unique for $M$ and $M\cong_{gr}N(\Sigma)$. Note that this defines an equivalence relation for $\Gamma$-graded modules. The next proposition helps to understand how gr-simple isoshift classes are. Note that, by Lemma 5.1, two gr-simple modules $M$ and $N$ are in the same isoshift class if and only if there exists $\sigma\in\Gamma$ such that $M\cong_{gr}N(\sigma)$.
@@
Let $S$ be a $\Gamma_{0}$-simple $R$-module, $D:=\operatorname{END}_{R}(S)$ and $\Gamma^{\prime}_{0}:=\Gamma^{\prime}_{0}(D)=\Gamma^{\prime}_{0}(S)$. The following assertions hold true.
@@
    There exists a bijection between $\Gamma^{\prime}_{0}/\sim$ and the isoshift classes of $\{S(e):e\in\Gamma^{\prime}_{0}\}$ that sends each $[e]\in\Gamma^{\prime}_{0}/\sim$ to the isoshift class of $S(e)$.
@@
for all $\gamma\in\Gamma$ such that $r(\gamma)=r(\sigma)$. We have that $\ker g^{\prime}$ is a proper graded submodule of $S(\sigma^{-1})$ and $\operatorname{im}g^{\prime}$ is a nonzero graded submodule of $S(r(\sigma))$. Since $S$ is $\Gamma_{0}$-simple, it follows that $S(\sigma^{-1})$ and $S(r(\sigma))$ are gr-simple $R$-modules. Therefore, $\ker g^{\prime}=0$ and $\operatorname{im}g^{\prime}=S(r(\sigma))$, i.e., $g^{\prime}$ is a gr-isomorphism. Let $h^{\prime}\in\operatorname{Hom_{gr-\mathit{R}}}(S(r(\sigma)),S(\sigma^{-1}))$ be the inverse of $g^{\prime}$. Extend $h^{\prime}$ to a $h\in\operatorname{Hom_{gr-\mathit{R}}}(S,S(\sigma^{-1}))$ defining $h(x)=0$ for all $x\in S(e)$ when $e\in\Gamma_{0}\setminus\{r(\sigma)\}$. Thus, $g\circ h:S\to S$ is the zero function on $S(e)$, for all $e\in\Gamma_{0}\setminus\{r(\sigma)\}$ and it is the identity on $S(r(\sigma))$, that is, $g\circ h=\mathds{1}_{r(\sigma)}$. On the other hand, $g$ (and therefore $h\circ g$) is the zero function on $S(e)$, for all $e\in\Gamma_{0}\setminus\{d(\sigma)\}$ and $h\circ g$ is the identity on $S(\sigma^{-1})$ ($=S(d(\sigma))$, as $R$-modules, by Lemma 2.6). Hence, $h\circ g=\mathds{1}_{d(\sigma)}$.
@@
If $S$ is a gr-simple $R$-module, then $D:=\operatorname{END}_{R}(S)$ is a gr-division ring with $\operatorname{supp}(D)\subseteq e\Gamma e$, where $e\in\Gamma_{0}$ is such that $S=S(e)$.∎
@@
Let $M$ be a $\Gamma$-graded $R$-module. Suppose $M=S_{1}\oplus\cdots\oplus S_{n}=T_{1}\oplus\cdots\oplus T_{m}$ where $n,m\in\mathbb{Z}_{>0}$ and $S_{1},...,S_{n},T_{1},...,T_{m}$ are gr-simple graded submodules of $M$. Then $n=m$ and there exists a permutation $\pi$ of $\{1,...,n\}$ such that $S_{i}\cong_{gr}T_{\pi(i)}$ for each $i=1,...,n$.
@@
Let $n>1$ and assume that the result is valid for $n-1$. Suppose $M=S_{1}\oplus\cdots\oplus S_{n}=T_{1}\oplus\cdots\oplus T_{m}$ as in the statement. Note that $m>1$. Given $1\leq i\leq n$ and $1\leq j\leq m$, let $p_{i}:M\to S_{i}$, $p^{\prime}_{j}:M\to T_{j}$ be the canonical projections and $\iota_{i}:S_{i}\to M$, $\iota^{\prime}_{j}:T_{j}\to M$ be the canonical inclusions. Then
@@
By the induction hypothesis, $n-1=m-1$ and there exists a bijection $\pi:\{1,...,k-1,k+1,...,n\}\to\{2,...,m\}$ such that $S_{i}\cong_{gr}T_{\pi(i)}$ for each $i=1,...,k-1,k+1,...,n$. Now just define $\pi(k)=1$. ∎
@@
Let $M$ be a $\Gamma$-graded $R$-module. Suppose $M=\bigoplus\limits_{i\in I}S_{i}=\bigoplus\limits_{j\in J}T_{j}$ where $S_{i},T_{j}$ are gr-simple graded submodules of $M$ for all $i\in I,j\in J$. Then $|I|=|J|$.∎
@@ 5.2. General results about gr-semisimple rings and modules
The first results of this section are basic facts on gr-semisimple modules. The following result will be used throughout and its proof can be found in [8, Lemma 51 and Propositions 52-53].
@@
[8, Proposition 57] Let $M$ be a $\Gamma$-graded $R$-module. If $M$ is a semisimple $R$-module, then $M$ is a gr-semisimple $R$-module.∎
In order to state the next result, we need a definition. We say that a $\Gamma$-graded $R$-module $Q$ is gr-injective if, for all gr-homomorphisms $j:M\to N$ and $g:M\to Q$ between $\Gamma$-graded $R$-modules with $j$ injective, there exists a gr-homomorphism $h:N\to Q$ such that $g=hj$ [8, Propositions 42 and 44].
@@
[8, Proposition 59] The following assertions are equivalent:
@@
Now we proceed in a similar way to [5, Section 4.6], [21, p. 35–36] and [10, Section 2]. Let $M$ be a $\Gamma$-graded $R$-module and $S$ be a gr-simple $R$-module. The isoshiftical component of type $S$ of $M$, denoted by $M_{S}$, is the sum of the graded submodules $T$ of $M$ that are in the same isoshift class of $S$. If $S^{\prime}$ is another gr-simple submodule of $M$ in the same isoshift class of $S$, then $M_{S}=M_{S^{\prime}}$. Thus $M_{S}$ depends only on the isoshift class of $S$. We denote by $\mathcal{S}(R)$ the set of isoshift classes of gr-simple $R$-modules. Thus, if $j\in\mathcal{S}(R)$ is the isoshift class of $S$, we can write $M_{j}$ instead of $M_{S}$.
@@
(2) By construction, $M_{j}\supseteq\bigoplus_{i\in I(j)}S_{i}$. Let $S$ be a gr-simple submodule of $M$ that belongs to $j$. Since $S$ is a cyclic $R$-module, $S\subseteq\bigoplus_{t=1}^{n}S_{i_{t}}$ for some $i_{1},\dotsc,i_{n}\in I$. Note that if the composition $S\stackrel{{\scriptstyle\iota}}{{\hookrightarrow}}\bigoplus_{t=1}^{n}S_{i_{t}}\stackrel{{\scriptstyle\pi_{i_{t}}}}{{\rightarrow}}S_{i_{t}}$ is not zero, where $\iota$ and $\pi_{i_{t}}$ are the natural inclusion and projection, respectively, then $S\cong_{gr}S_{i_{t}}$. Thus, $i_{1},\dotsc,i_{n}$ can be chosen to belong to $I(j)$ Hence $S\subseteq\sum_{i\in I(j)}S_{i}$. Therefore, $M_{j}\subseteq\bigoplus_{i\in I(j)}S_{i}$. ∎
@@
:=$\mathcal{F}:=\{T\textrm{ graded right ideal of $R$}:\textrm{there exists $\sigma\in e\Gamma$ such that $T_{R}\cong_{gr}S(\sigma)$}\}.$
@@
Let $\{R_{i}:i\in I\}$ be a infinite family of rings with unity. Each $R_{i}$ is a $I\times I$-graded ring via $R_{i}=(R_{i})_{(i,i)}$ and we have
@@
Now we are ready to show the main result of this subsection. We follow very close the proof of the ungraded case given in [21, p. 35-36].

### v2

@@ 4.2. Graded modules over gr-division rings
The proofs of Theorem 4.4 and Corollary 4.6 follow very much the pattern for group graded division rings [20, Section 1.4].
@@
Let $D$ be a $\Gamma$-graded division ring and $M$ be a $\Gamma$-graded $D$-module. The following assertions hold:
@@
$\mathcal{F}=\left\{(x_{i})_{i\in Q}\in\prod\limits_{i\in Q}M_{\gamma_{i}}\mid I\subseteq Q,\,(x_{i})_{i\in Q}\textrm{ is pseudo-linearly independent}\right\}.$
@@
If $D$ is a $\Gamma$-graded division ring, then every $\Gamma$-graded $D$-module has a pseudo-basis by Theorem 4.4(1). Unlike in the group graded context [7, Theorem 3.3], the converse is not true. We will deal with this situation in Section 7.
@@
Let $D$ be a $\Gamma$-graded division ring and $M$, NN be $\Gamma$-graded $D$-modules. For all $(\gamma_{i})_{i\in I}\in\Gamma^{I}$, pseudo-linear independent sequence $(x_{i})_{i\in I}\in\prod\limits_{i\in I}M_{\gamma_{i}}$, $\sigma\in\Gamma$ and $(y_{i})_{i\in I}\in\prod\limits_{i\in I}N_{\sigma\gamma_{i}}$, there exists $g\in\operatorname{HOM}_{D}(M,N)_{\sigma}$ such that $g(x_{i})=y_{i}$ for each $i\in I$.
@@ 4.3. Graded linear algebra over gr-division rings
$\operatorname{M}_{m\times n}(R)[\overline{\alpha}][\overline{\beta}]=\left\{\begin{pmatrix}a_{11}&\cdots&a_{1n}\\ \vdots&\ddots&\vdots\\ a_{m1}&\cdots&a_{mn}\end{pmatrix}\in\operatorname{M}_{m\times n}(R)\ \bigg|\ a_{ij}\in R_{\alpha_{i}\beta_{j}^{-1}}\right\},$
where we follow the convention $R_{\alpha_{i}\beta_{j}^{-1}}=\{0\}$ if $d(\alpha_{i})\neq r(\beta_{j}^{-1})=d(\beta_{j})$.
@@
Define $f]_{\mathcal{B}\mathcal{C}}\in\operatorname{M}_{m\times n}(R)[\overline{\alpha}][\overline{\beta}]$ as the matrix
$f]_{\mathcal{BC}}=\Big[(f(u_{1}))_{\mathcal{C}}\ (f(u_{2}))_{\mathcal{C}}\ \dotsb(f(u_{n}))_{\mathcal{C}}\Big]=\begin{pmatrix}a_{11}&a_{12}&\ldots&a_{1n}\\ a_{21}&a_{22}&\ldots&a_{2n}\\ \vdots&\ddots&\ddots&\vdots\\ a_{m1}&a_{m2}&\ldots&a_{mn}\end{pmatrix}$.
@@
$f]_{\mathcal{BC}}\cdot(x)_{\mathcal{B}}=(f(x))_{\mathcal{C}}\ \textrm{ for all }x\in\operatorname{h}(M)$ (4.1)
and that $f]_{\mathcal{BC}}$ is the unique matrix in $\operatorname{M}_{m\times n}(R)[\overline{\alpha}][\overline{\beta}]$ that satisfies (4.1).
@@
$f]_{\mathcal{BC}}\cdot[g]_{\mathcal{AB}}=[f\circ g]_{\mathcal{AC}}.$
@@
The following result, shows that all for ranks just defined are equal over a graded division ring. We follow very close the proof in [26, Exercises 13.13, 13.14] where the result is proved for (ungraded) division rings.
@@ 5.1. Gr-simple modules and Schur’s lemma
When considering rings and modules without a grading, it is useful to study simple modules up to isomorphism. In the graded context, one has to take into account the shifts of the simple modules. Inspired by the idea of [13, p. 395], we say that two $\Gamma$-graded $R$-modules $M$ and $N$ are in the same isoshift class if there exists $\Sigma\subseteq\Gamma$ such that $\Sigma$ is fully $r$-unique for $N$, $\Sigma^{-1}$ is fully $r$-unique for $M$ and $M\cong_{gr}N(\Sigma)$. Note that this defines an equivalence relation for $\Gamma$-graded modules. The next proposition helps to understand how gr-simple isoshift classes are. Note that, by Lemma 5.1, two gr-simple modules $M$ and $N$ are in the same isoshift class if and only if there exists $\sigma\in\Gamma$ such that $M\cong_{gr}N(\sigma)$.
@@
Let $S$ be a $\Gamma_{0}$-simple $R$-module, $D:=\operatorname{END}_{R}(S)$ and $\Gamma^{\prime}_{0}:=\Gamma^{\prime}_{0}(D)=\Gamma^{\prime}_{0}(S)$. The following assertions hold:
@@
    There exists a bijection between $\Gamma^{\prime}_{0}/\sim$ and the isoshift classes of $\{S(e):e\in\Gamma^{\prime}_{0}\}$ that sends each $e]\in\Gamma^{\prime}_{0}/\sim$ to the isoshift class of $S(e)$.
@@
for all $\gamma\in\Gamma$ such that $r(\gamma)=r(\sigma)$. We have that $\ker g^{\prime}$ is a proper graded submodule of $S(\sigma^{-1})$ and $\operatorname{im}g^{\prime}$ is a nonzero graded submodule of $S(r(\sigma))$. Since $S$ is $\Gamma_{0}$-simple, it follows that $S(\sigma^{-1})$ and $S(r(\sigma))$ are gr-simple $R$-modules. Therefore, $\ker g^{\prime}=0$ and $\operatorname{im}g^{\prime}=S(r(\sigma))$, i.e., $g^{\prime}$ is a gr-isomorphism. Let $h^{\prime}\in\operatorname{Hom_{gr-\mathit{R}}}(S(r(\sigma)),S(\sigma^{-1}))$ be the inverse of $g^{\prime}$. Extend $h^{\prime}$ to a $h\in\operatorname{Hom_{gr-\mathit{R}}}(S,S(\sigma^{-1}))$ defining $h(x)=0$ for all $x\in S(e)$ when $e\in\Gamma_{0}\setminus\{r(\sigma)\}$. Thus, $g\circ h:S\to S$ is the zero function on $S(e)$, for all $e\in\Gamma_{0}\setminus\{r(\sigma)\}$ and it is the identity on $S(r(\sigma))$, that is, $g\circ h=\mathds{1}_{r(\sigma)}$. On the other hand, $g$ (and therefore $h\circ g$) is the zero function on $S(e)$, for all $e\in\Gamma_{0}\setminus\{d(\sigma)\}$ and $h\circ g$ is the identity on $S(\sigma^{-1})$ ($S(d(\sigma))$, as $R$-modules, by Lemma 2.6). Hence, $h\circ g=\mathds{1}_{d(\sigma)}$.
@@
If $S$ is a gr-simple $R$-module, then $D:=\operatorname{END}_{R}(S)$ is a gr-division ring with $\operatorname{supp}(D)\subseteq e\Gamma e$, where $e\in\Gamma_{0}$ is such that S=S(e)S=S(e).∎
@@
Let $M$ be a $\Gamma$-graded $R$-module. Suppose $M=S_{1}\oplus\cdots\oplus S_{n}=T_{1}\oplus\cdots\oplus T_{m}$ where $n,m\in\mathbb{Z}_{>0}$ and $S_{1},...,S_{n},T_{1},...,T_{m}$ are gr-simple graded submodules of $M$. Then $n=m$ and there exists a permutation $\pi$ of $\{1,...,n\}$ such that $S_{i}\cong_{gr}T_{\pi(i)}$ for each i=1,,ni=1,...,n.
@@
Let $n>1$ and assume that the result is valid for n1n-1. Suppose $M=S_{1}\oplus\cdots\oplus S_{n}=T_{1}\oplus\cdots\oplus T_{m}$ as in the statement. Note that $m>1$. Given $1\leq i\leq n$ and $1\leq j\leq m$, let $p_{i}:M\to S_{i}$, $p^{\prime}_{j}:M\to T_{j}$ be the canonical projections and $\iota_{i}:S_{i}\to M$, $\iota^{\prime}_{j}:T_{j}\to M$ be the canonical inclusions. Then
@@
By the induction hypothesis, n1=m1n-1=m-1 and there exists a bijection $\pi:\{1,...,k-1,k+1,...,n\}\to\{2,...,m\}$ such that $S_{i}\cong_{gr}T_{\pi(i)}$ for each i=1,,k1,k+1,,ni=1,...,k-1,k+1,...,n. Now just define $\pi(k)=1$. ∎
@@
Let $M$ be a $\Gamma$-graded $R$-module. Suppose $M=\bigoplus\limits_{i\in I}S_{i}=\bigoplus\limits_{j\in J}T_{j}$ where $S_{i},T_{j}$ are gr-simple graded submodules of $M$ for all $i\in I,j\in J$. Then |I|=|J||I|=|J|.∎
@@ 5.2. General results about gr-semisimple rings and modules
The first results of this section are basic facts on gr-semisimple modules. The following result will be used throughout the paper and its proof can be found in [11, Lemma 51 and Propositions 52-53].
@@
[11, Proposition 57] Let $M$ be a $\Gamma$-graded $R$-module. If $M$ is a semisimple $R$-module, then $M$ is a gr-semisimple $R$-module.∎
In order to state the next result, we need a definition. We say that a $\Gamma$-graded $R$-module $Q$ is gr-injective if, for all gr-homomorphisms $j:M\to N$ and $g:M\to Q$ between $\Gamma$-graded $R$-modules with $j$ injective, there exists a gr-homomorphism $h:N\to Q$ such that $g=hj$ [11, Propositions 42 and 44].
@@
[11, Proposition 59] The following assertions are equivalent:
@@
Now we proceed in a similar way to [8, Section 4.6], [24, p. 35–36] and [13, Section 2]. Let $M$ be a $\Gamma$-graded $R$-module and $S$ be a gr-simple $R$-module. The isoshiftical component of type $S$ of $M$, denoted by $M_{S}$, is the sum of the graded submodules $T$ of $M$ that are in the same isoshift class of $S$. If $S^{\prime}$ is another gr-simple submodule of $M$ in the same isoshift class of $S$, then $M_{S}=M_{S^{\prime}}$. Thus $M_{S}$ depends only on the isoshift class of $S$. We denote by $\mathcal{S}(R)$ the set of isoshift classes of gr-simple $R$-modules. Thus, if $j\in\mathcal{S}(R)$ is the isoshift class of $S$, we can write $M_{j}$ instead of $M_{S}$.
@@
(2) By construction, $M_{j}\supseteq\bigoplus_{i\in I(j)}S_{i}$. Let $S$ be a gr-simple submodule of $M$ that belongs to $j$. Since $S$ is a cyclic $R$-module, $S\subseteq\bigoplus_{t=1}^{n}S_{i_{t}}$ for some $i_{1},\dotsc,i_{n}\in I$. Note that if the composition $S\mathrel{{\mathop{\hookrightarrow}\limits^{\iota}}}\bigoplus_{t=1}^{n}S_{i_{t}}\mathrel{{\mathop{\rightarrow}\limits^{p_{i_{t}}}}}S_{i_{t}}$ is not zero, where $\iota$ and $p_{i_{t}}$ are the natural inclusion and projection, respectively, then $S\cong_{gr}S_{i_{t}}$. Thus, $i_{1},\dotsc,i_{n}$ can be chosen to belong to $I(j)$ . Hence $S\subseteq\sum_{i\in I(j)}S_{i}$. Therefore, $M_{j}\subseteq\bigoplus_{i\in I(j)}S_{i}$. ∎
@@
[…]
@@
Let $\{R_{i}:i\in I\}$ be an infinite family of rings with unity. Each $R_{i}$ is a $I\times I$-graded ring via $R_{i}=(R_{i})_{(i,i)}$ and we have
@@
Now we are ready to show the main result of this subsection. We follow very close the proof of the ungraded case given in [24, p. 35-36].